Circular No. 7750

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001fv IN NGC 3512 G. M. Hurst, Basingstoke, England, reports the discovery by M. Arm-

strong, Rolvenden, England, of an apparent supernova (mag 16.4) on an unfiltered CCD image taken on Nov. 3.175 UT in the course of the U.K. Nova/Supernova Patrol; the object was at mag 16.8 on images obtained on Nov. 10.16. SN 2001fv is located at α = 11^{h}04^{m}01^{s}.66, δ = +28^{o}01′55′′.7, which is 21′′ west and 10′′.5 south of the center of NGC 3512. The new object is not present Armstrong’s image from 2000 Dec. 22 (limiting mag 19.0) or on Palomar Sky Survey images taken on 1955 Mar. 22, 1992 May 2 (blue, limiting mag 22.5), and 1998 May 21 (red, limiting mag 20.5)

CI AQUILAE B. E. Schaefer, University of Texas, Austin, writes that an outburst of

the recurrent nova CI Aql has been found on 12 Harvard College Observatory archival photographs spanning 1941 May 2–1942 Aug. 4, yielding the following representative B magnitudes: 1941 May 2, 13.8; 24, 14.7; Sept. 11, 14.9; 1942 May 18, 16.5; Aug. 4, 16.65. The peak brightness may well have been brighter, with few useful plates in the six months prior to this span. This outburst’s light curve is similar to that in 2000. Deep plates show that the normal magnitude range for this eclipsing binary is 17.0– 17.8. Schaefer suggests that the 1917 discovery could indicate a recurrence timescale of ∼ 20 yr, with missed outbursts around 1960 and 1980.

SUNGRAZING COMETS Further to IAUC 7746, D. Hammer reports his measurements of addi-

tional sungrazing comets found on SOHO website images by X. M. Zhou (C/2001 T5), R. Kracht (C/2001 T6), M. Oates (C/2001 T7), and S. Hönig (C/2001 U7). All the objects were seen in the C2 coronagraph except for C/2001 T5 (the only non-Kreutz object, seen only in C3); C/2001 T7 and C/2001 U7 were also seen in C3. The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  2001 UT  α_{2000}  δ_{2000}  MPEC
C/2001 T5 Oct. 14.863  13^{h}17.^{m}3  − 3^{o}45′  2001-V10
C/2001 T6  15.535  13 14.0  − 9 25  2001-V10
C/2001 T7  6.513  12 33.0  − 6 07  2001-V19
C/2001 U7  18.513  13 07.1  −12 16  2001-V31
2001 November 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green